# Improving Genetic Programming for Symbolic Regression with Equality Graphs

### v1

@@ 1. Introduction
Symbolic regression (SR) searches for a mathematical function that approximates a set of data points. It can be used to find nonlinear regression models (Koza, 1992; Kronberger et al., 2024a) or for scientific discovery (Kronberger et al., 2024a; Cranmer, 2023; Udrescu and Tegmark, 2020; Russeil et al., 2024; Cao et al., [n. d.]; Lamb and Harrington, [n. d.]).
The current state-of-the-art SR algorithms (La Cava et al., 2021; de Franca et al., 2024) use genetic programming (GP) as the main search engine. One of such improvements, used by several state-of-the-art implementations, is the search for a parametric function that replaces the constants with adjustable parameters that are fitted using nonlinear optimization techniques.
Not only is the search for a symbolic model NP-hard (Virgolin and Pissis, [n. d.]) but when searching for a parametric model, it usually requires the solution to a multimodal optimization problem, which by itself is NP-hard (Murty and Kabadi, 1985) and can hinder the search for the optimal solution.
@@
To make matters worse, the usual way of encoding mathematical expressions as symbolic expression trees, allows GP to visit different but equivalent expressions (Kronberger et al., 2024b) that evaluate to the same values. These equivalent expressions may be unnecessarily large and contain redundant parameters, which reduces the probability of finding their optimal values (de Franca and Kronberger, 2023; Kronberger and de França, 2024). Even for the simple expression we can produce an infinite number of equivalent expressions considering that are fitting parameters, for example , , , etc. are all different parameterizations of the same expression.
@@
GP cannot differentiate between equivalent forms of a given expression, and simplification heuristics are often insufficient, as seen in (de Franca and Kronberger, 2023). Some authors (Keller and Banzhaf, 1999; Miller and Smith, 2006) argue that redundancy is necessary to allow the algorithm to navigate through the search space, as these equivalent expressions are guaranteed to have the same accuracy, allowing the search to keep multiple genetically different variations of solution candidates in the hopes of finding a better solutions. However, we do not know what would happen to GP search dynamics if we try to prevent keeping semantic duplicates in the GP population.
@@
Equality saturation (Willsey et al., 2021) uses a set of equivalence rules and produces all equivalent forms of a given expression. It alleviates the phase ordering problem in the optimization phase when compiling computer programs. Given a program represented as a symbolic expression tree, and a set of equivalence rules, it iteratively applies the rules and stores all equivalent programs in a compact data structure called equality graph (e-graph). The main idea is that upon saturation, the graph contains all equivalent forms of the original program and the optimal form can be extracted from the e-graph using a heuristic. This technique was previously used in the context of SR in (de Franca and Kronberger, 2023; Kronberger and de França, 2024) to investigate the problem of overparameterization that can negatively affect the fitting of numerical parameters. The e-graph has another interesting feature that can be exploited by SR algorithms: it contains a database of all visited patterns and their equivalent forms that can be easily matched against new candidate expressions.
In this work, we investigate the benefits of exploiting the e-graph to increase the chances of generating a novel expression when applying crossover and mutation. In short, the proposed crossover operator will limit the selection of the second parent to a subset of subtrees that can only generate an unvisited expression. For the subtree mutation, it will replace a random subtree from the original tree with a random subtree generating the new tree . If has already been visited, the root of will be changed to a random choice of nonterminals that generates an unvisited expression.
@@
These operators are tested inside a simple GP implementation based on tinyGP (Sipper, 2019) and compared against tinyGP with parameter fitting and two state-of-the-art algorithms: Operon (Burlacu et al., 2020b) and PySR (Cranmer, 2023). The results show that this simple111We acknowledge that the modification is simple provided we have a working implementation of e-graph. modification can improve the performance of tinyGP to an extent that it becomes competitive (and in some aspects better) than the state-of-the-art. The use of an e-graph as a support structure for GP brings new light to symbolic regression and GP with the possibility of exploring the accumulated history of the search process and even combining the history of multiple searches.
This paper is organized such that in Section 2 we will summarize the related works in symbolic regression. Section 3 will explain the basic concepts of equality saturation and the e-graph data structure. In Section 4 we will detail the proposed modification to simple GP. Section 5 will show the experiment methods used in this paper followed by Section 6 where we report and discuss the results. Finally, Section 7 gives some final remarks and expectations for the future.
@@ 2. Related work
Kronberger et al. (Kronberger et al., 2024b) studied the inefficiency of a simple GP comparing with the enumerated search space (Bartlett et al., 2023) and using equality saturation to count the percentage of unique expressions generated during the GP search. They found that from the total of visited expressions during the search, only around were unique. This not only wastes computational resource but it also shows that, at some point, GP fails to explore different regions of the search space.
Many authors observed improvements in the obtained solutions when applying any form of simplification during the search (Cao et al., 2023; Randall et al., 2022; Rivero et al., 2022; Seidyo Imai Aldeia et al., 2024) and, as a side effect, it stimulates the diversity of the population (Burlacu et al., 2019, 2020a).
Equality saturation has been used in the context of symbolic regression as a support tool to study the behavior of the search. Many state-of-the-art SR algorithms have a bias towards creating expressions with redundant numerical parameters (de Franca and Kronberger, 2023; Kronberger and de França, 2024). This redundancy can increase the chance of failing to correctly optimize such parameters, leading to sub-optimal solutions. In (Kronberger et al., 2024b) this technique was used to detect the equivalent expressions visited during the GP search. So far, the equality saturation technique was not used during the GP search to improve the quality of the solutions.
@@ 3. Equality saturation and e-graphs
Equality saturation (Tate et al., 2009) was proposed as a solution to the phase ordering problem in compiler optimization. This problem occurs when optimizing a program by applying a set of rewrite rules sequentially while dropping the information about the previous versions of the program. If the optimization follows a non-optimal sequence, it will lead to a sub-optimal program.
Equality saturation solves this issue by applying all of the optimization rules in parallel while keeping the intermediate transformations in a compact form using the data structure called e-graph.
@@
Fig. 1a illustrates an example of an e-graph. Each solid box represents an e-node that contains a symbol of the expression. The dashed boxes group a set of e-nodes together and it is called an e-class, each e-class is assigned an e-class id (number in the bottom right of an e-class box). The main property of an e-class is that, no matter which e-node is chosen during the traversal, it will lead to an equivalent expression with all other e-nodes.
@@
We can see that in the middle box (e-class id ), if we follow through it will generate the expression and if we follow through it will generate . The abstract description of the algorithm is very simple, though a concrete and optimal implementation requires the use of advanced techniques and data structures. The main idea is: i) match all the equivalence rules in the current state of the e-graph, ii) apply the rules creating new e-classes, iii) merge the equivalent e-classes, iv) repeat until saturation (i.e., no changes occur).
@@
The current state-of-the-art implementation (Willsey et al., 2021)222https://docs.rs/egg/latest/egg/, called egg333egg stands for e-graphs good, encodes the e-graph as a map of an e-node to an e-class, a map of the e-class ids to their structure, and a set of e-classes that still needs analysis to maintain consistency. The structure of an e-class contains the information about the e-nodes it contains, a list of the parent e-nodes, and additional information also referred to as semantic analysis. This implementation also maintains a database of patterns that allows the algorithm to efficiently match patterns inside the e-graph structure.
@@
Usually, the e-graph is used to represent a single program or expression and their equivalent forms. But, the structure can keep any number of expressions as long as we keep a list of the e-classes ids representing the root of each expression. For example, if we insert the expressions into the e-graph, we would end up with Fig. 1a, minus the equivalent relations. As a result, we would keep the list representing the ids of the expressions we inserted. New expressions are added bottom-up. Starting from a terminal, the algorithm checks whether it already exists in the e-graph and returns its e-class id, if it does not exist, it creates a new id. When adding an internal node, the algorithm first converts it to an e-node by replacing its children by their e-class ids and then it checks whether it already exists in the e-graph, returning the corresponding e-class id or a new one. In our example from Fig. 1, if we try to add the expression , it would first retrieve the e-class ids for the terminals and , then it would return the e-class id corresponding to the e-node , where underlined numbers correspond to e-class ids. Finally, it would create a new e-class with id and the e-node . This mechanism allows us to compactly store a set of expressions and readily assert whether an expression already exists in the structure.
In our implementation, we made a series of improvements to this data structure as we will describe in the next section.
@@
The proposed algorithm, eggp (e-graph genetic programming), follows the same structure as the traditional GP. The initial population is created using ramped half-and-half respecting a maximum size and maximum depth parameter (Koza, 1992) and, for a number of generations, it will choose two parents using tournament selection, apply the subtree crossover with probability followed by the subtree mutation with probability , when the offsprings replace the current population following a dominance criteria.
The key differences of eggp are:
new solutions are inserted into the e-graph followed by one step of equality saturation to find and store some of the equivalent expressions of the new offspring.
the current population is replaced by the set of individuals formed by: the Pareto front, the next front after excluding the first Pareto-front, and a selection of the last offspring at random until it reaches the desired population size.
@@
Apart from that, we do not make use of other mutation operators commonly used in the literature (Burlacu et al., 2020b; Cranmer, 2023) or advanced techniques to stimulate diversity such as island models. Notice that a single step of equality saturation will not guarantee the insertion of all equivalent expressions in the e-graph but, as we apply more iterations, the e-graph can grow exponentially large. This issue is amplified by the fact that we are storing multiple expressions. As we will see in Sec. 6, the single step seems to be sufficient to improve the results and the benefits of increasing the number of steps is a subject for a future research. Regarding the Pareto-front extracted from the entire history, this is equivalent to the traditional NSGA-II algorithm (Deb et al., 2000) as the dominance relation is transitive. Keeping two ranks of dominance and filling up the remainder of the population with new expressions is meant to stimulate the combination of new expressions (exploration) while keeping the best fronts (exploitation). As a side-note, in this implementation we keep a database of generated expressions sorted by fitness and size (both objectives used in this work), so the Pareto-front can be retrieved in where is the number of extracted individuals. A new expression can be inserted into this structure in , where is the number of evaluated expressions so far. Finally, if we add the expression , equality saturation will generate the equivalent form (constants are replaced by parameters), this expression is stored in the database of expressions with size .
4.1. Crossover and mutation
Since the algorithm keeps the whole history of visited expression and its equivalent forms, we can exploit this information to increase the probability of generating unvisited expressions in the perturbation operators. For the crossover (Fig. 2a), the main idea is that it first chooses a crossover point from the first parent and then it randomly chooses a subtree of the second parent which generates an unvisited expression upon combination.
For the mutation operator (Fig. 2b), it chooses a point of the tree to generate a new subtree and, after generating the new subtree, it checks whether it generates an unvisited expression. With a negative answer, it randomly changes the root node of this subtree from a choice of all the possible non-terminals that can ensure novelty. In the event of impossibility of generating a new expression, a random recombination is returned.
To verify whether the combination of the original tree with the subtree replacing the -th node will generate an unvisited expression, we traverse the tree in pre-order traversal until we reach the node we want to replace (-th node). While traversing the tree, we build a list of -functions equivalent to a function that, given a tree , it creates a new tree with the operator of as the root, replacing one child with and keeping the other child of (if it is a binary operator). So, if the next step in the traversal is to the left, this function will build a tree rooted at with the left child being replaced by and the right child as , this is illustrated in Alg. 1. Upon reaching the desired node, we traverse the tree upwards by checking whether the current node exists in the e-graph and, if the assertion is false, it will apply the first function of the list to , generating the parent e-node, and calling this same function recursively (Alg. 2). Given that none of the ancestors exist in the e-graph, the algorithm confirms that this is an unvisited expression.
@@ 4.1. Crossover and mutation
The full implementation of eggp has nine hyperparameters: number of generations, population size, maximum expression size, likelihood function (MSE, Gaussian, Poisson, Bernoulli, ROXY (Lelli et al., 2017)), number of iterations and retries for the parameter optimization, probabilities of crossover and mutation, and the list of non-terminals.
The algorithm is implemented in Haskell using the srtree444https://github.com/folivetti/srtree library for symbolic regression and with an adapted implementation of equality saturation based on hegg555https://github.com/alt-romes/hegg. The binaries are available at https://github.com/folivetti/srtree/releases/tag/v2.0.1.0.
@@ 5. Experiments
Operon is a carefully crafted implementation of GP for symbolic regression with runtime performance in mind and a good set of default hyperparameters. It incorporates multiple mutation operators that allow a finer perturbation of a solution. Besides, it envelops every variable node with a scaling parameter adjusted using nonlinear optimization. PySR also supports the same mutation operators and nonlinear optimization of the parameters, it stands out with the use of an island model capable of keeping the diversity of the population to stimulate the exploration of the search space. It also applies a simplification heuristic on a selection of the expressions. Both PySR and Operon uses multi-objective optimization optimizing accuracy and size as default. The research questions we want to address with the experiments are:
What is the impact of increasing the probability of generating novelty, when compared to a minimalistic implementation such as tinyGP?
How close does eggp get to the state-of-the-art without resorting to more advanced concepts such as specialized mutation operators, enforcing the placement of numerical parameters, and island model?
@@
We have fixed all the common hyperparameters to a standard value and performed a hyperparameter tuning with two different settings for each algorithm varying only whether to have a larger population with fewer iterations or a smaller population with more iterations. The only differences in settings are: for PySR we are using populations in its island model, so the size of each population is of the population size for the other algorithms, earlier experiments with the first set of benchmarks revealed that PySR performs significantly worse when using a single population; for Operon, we perform a maximum of optimization iterations instead of iterations with different starting points since it does not support multiple restarts for the parameter optimization; for eggp, of the training data is separated and used as a validation set to calculate the fitness, while the parameters are fitted using the remaining . Notice that tinyGP is the only single-objective approach and will be left out of the hypervolume comparison.
@@
| number of evaluations | |
@@
For the first set of experiments, we picked four datasets from the original SRBench (de Franca et al., 2024) and split each one into training and test sets with a ratio of . We executed each configuration times and stored the test errors of each run. Once we have the optimal hyperparameters, we evaluate these algorithms in the benchmarks of the current version of SRBench666https://github.com/cavalab/srbench/discussions/174#discussioncomment-10285133 for the reduced SRBench track. This set is supposed to be representative as it contains datasets with different characteristics. For this experiment, we applied a -fold cross-validation repeating the experiment times, creating a total of runs. Finally, we picked some real-world datasets from the literature corresponding to data from different fields, these data already have pre-determined training and test sets, as such we will run repetitions of each experiment. For every experiment, we store the final Pareto front and will report the performance plot, the AUC and average rank among all datasets and statistical test of the ranks, the average and standard deviation of the running time, the hypervolume of the Pareto front. All algorithms were restricted to a single core to ensure equal conditions.
Table 2 shows the datasets of each set of experiment with their corresponding number of data points, features and the chosen maximum size parameter. 777The maximum size for Operon is set to of the maximum size because, internally, Operon will not count the scale coefficients of a terminal towards the model size. This factor enforces Operon to search on a similar search space as the other algorithms.
@@
| hyperparameter tuning | |||
| 523_analcatdata_neavote | |||
| 527_analcatdata_election2000 | |||
| 635_fri_c0_250_10 | |||
| 1029_LEV | |||

### v2

@@ 1. Introduction
Symbolic regression (SR) (Koza, 1992; Kronberger et al., 2024a) searches for a mathematical function that approximates a set of data points often used for scientific discovery (Schmidt and Lipson, 2009; Kronberger et al., 2024a; Cranmer, 2023; Udrescu and Tegmark, 2020; Russeil et al., 2024; Cao et al., 2024). The current SotA (La Cava et al., 2021; de Franca et al., 2024) uses genetic programming (GP) as the main search engine incorporating numerical parameters that can be fitted into the data using optimization techniques. The search for a symbolic model is NP-hard (Virgolin and Pissis, 2022) and when searching for a parametric model, it also requires the solution to a multimodal optimization problem, which by itself is NP-hard (Murty and Kabadi, 1985) and can hinder the search for the optimal solution.
@@
To make matters worse, the usual way of encoding mathematical expressions as symbolic expression trees, allows GP to visit semantically equivalent expressions111in this paper, we will refer to semantically equivalence as simply equivalence. with different syntax (Kronberger et al., 2024b). These equivalent expressions may be unnecessarily large and with redundant parameters, reducing the probability of finding their optimal values (de Franca and Kronberger, 2023; Kronberger and Olivetti de França, 2025). Even for the simple expression we can produce an infinite number of equivalent expressions considering that are fitting parameters, for example , , , are all different parameterizations of the same expression.
@@
GP cannot easily differentiate between equivalent expressions, and applying simplification heuristics are often insufficient, as seen in (de Franca and Kronberger, 2023). Some authors (Keller and Banzhaf, 1999; Miller and Smith, 2006) argue that redundancy is necessary to allow the algorithm to navigate through the search space, as these equivalent expressions are guaranteed to have the same accuracy, allowing the search to keep multiple genetically different variations of solution candidates in the hopes of finding better solutions. However, this supposition could not be validate as there were no means to efficiently verify equivalence.
@@
Equality saturation (Willsey et al., 2021) can produce all the equivalent expressions of a given expression through the parallel application of a set of equivalence rules. Given an expression represented as a directed acyclic graph, and a set of equivalence rules, it iteratively applies the rules and stores all equivalent programs in a compact data structure called equality graph (e-graph). The main idea is that upon saturation, the graph will contain all equivalent forms of the original program and the optimal form can be extracted from the e-graph using a heuristic cost function. This technique was previously used in the context of SR in (de Franca and Kronberger, 2023; Kronberger and Olivetti de França, 2025) to investigate the problem of overparameterization that can negatively affect the fitting of numerical parameters. The e-graph has another feature that can be exploited by SR algorithms: it implements an efficient pattern matching algorithm that can answer whether a given expression, or any of its equivalents, are already stored in the e-graph structure.
In this work we introduce eggp, a GP algorithm that exploits the pattern matching capabilities of the e-graph to try to enforce the generation of unvisited expressions when applying the crossover and mutation operators. In this context, unvisited expressions mean any expression, or their equivalents, that was not previously evaluated during the history of search. In short, after choosing the crossover point of the first parent, the choices of points of the second parent are limited to those that will ensure the generation of an unvisited expression. For the mutation, after choosing a node of the expression at random, it will generate a new subtree, limiting the choice of its root node to the set that will ensure the generation of an unvisited expression. This procedure will enforce the introduction of novel solutions not only w.r.t. the current population but to the entire history of the search.
The research questions we want to address in this paper are:
What is the impact of increasing the probability of generating novelty, when compared to a minimalist implementation of GP for SR?
How close does eggp get to the state-of-the-art without resorting to more advanced concepts such as specialized mutation operators, enforcing the placement of numerical parameters, and promoting diversity throug island model?
@@
These operators are tested inside a minimalist GP implementation, and compared against this same algorithm with the original subtree operators, and two high performant algorithms: Operon (Burlacu et al., 2020b) and PySR (Cranmer, 2023). The results show that this simple modification, provided we have a working implementation of the e-graph, can improve the performance of this minimalist GP to an extent that it becomes competitive (and in some aspects better) than the state-of-the-art. The use of an e-graph as a support structure for GP brings new light to symbolic regression and GP with the possibility of exploring the accumulated history of the search process and even combining the history of multiple searches. This paper is organized such that in Section 2 we will summarize the related works in symbolic regression. Section 3 will explain the basic concepts of equality saturation and the e-graph data structure. In Section 4 we will detail the proposed modifications to the subtree operators. Section 5 and 6 will respectively detail the experiment methods, report and discuss the results. Finally, Section 7 will provide some final remarks and expectations for the future.
@@ 2. Related work
Kronberger et al. (Kronberger et al., 2024b) studied the inefficiency of a simple GP comparing with the enumerated search space (Bartlett et al., 2024) and using equality saturation to count the percentage of unique expressions generated during the GP search. They found that from the total of visited expressions during the search, only around were unique. This not only wastes computational resources but it also shows that, at some point, GP fails to explore different regions of the search space. Many authors observed improvements in the obtained solutions when applying any form of simplification during the search (Cao et al., 2023; Randall et al., 2022; Rivero et al., 2022; Imai Aldeia et al., 2024) while also stimulating the diversity of the population (Burlacu et al., 2019, 2020a).
Equality saturation has been used in the context of symbolic regression as a support tool to study the behavior of the search. Many state-of-the-art SR algorithms have a bias towards creating expressions with redundant numerical parameters (de Franca and Kronberger, 2023; Kronberger and Olivetti de França, 2025). This redundancy can increase the chance of failing to correctly optimize such parameters, leading to sub-optimal solutions. In (Kronberger et al., 2024b) this technique was used to detect the equivalent expressions visited during the GP search. So far, the equality saturation technique was not used during the GP search to improve the quality of the solutions.
Semantic similarity is frequently studied in the GP literature, either to improve the population diversity, the locality of the perturbation operators, or to understand the dynamics of GP search (Vanneschi et al., 2014). The semantic aware operators (Uy et al., 2010) introduce locality by replacing subtrees of an expression with semantically similar trees. Another approach is to combine two expressions by generating a random expression with a codomain in the range and creating the combined expression , ensuring a balance between the semantics of and (Moraglio et al., 2012). In (Ruberto et al., 2019), the authors introduce the equivalence function that determines whether two expressions are equivalent if the differences in their behavior in the semantic space is constant. Using this idea, they implement a filtering mechanism that rejects any offspring that is equivalent to any expression in the current population.
Given the definition of mathematical equivalence, stating that , where is the variable domain, it is important to highlight that in the Semantic GP literature, semantic equivalence is often calculated using a limited number of data points (), which cannot guarantee equivalence, but is sufficient for an approximate measure of semantic similarity. Using equality saturation we can produce and store equivalent expressions without the need of evaluation, as explained in the next section. In this paper, we are concerned with semantic equality and the means to enforce the creation of unvisited expressions, regardless of locality.
@@ 3. Equality saturation and e-graphs
Equality saturation (Tate et al., 2009) was proposed as a solution to the phase ordering problem in compiler optimization. This problem occurs when optimizing a program by applying a set of rewrite rules sequentially while dropping the information about the previous versions of the program. If the optimization follows a non-optimal sequence, it will lead to a sub-optimal program. Equality saturation solves this issue by applying all of the optimization rules in parallel while keeping the intermediate transformations in a compact form using the data structure called e-graph.
@@
Fig. 1a illustrates an example of an e-graph. Each solid box represents an e-node that contains a symbol of the expression. The dashed boxes, called e-class, group a set of e-nodes together. Each one of these e-classes is assigned an id (number in the bottom right of an e-class box). The main property of an e-class is that, no matter which e-node is chosen during the traversal, it will lead to an equivalent expression to all other e-nodes of the same e-class.
@@
Looking at the middle box (e-class id ), if we follow through it will generate the expression and if we follow through it will generate . The abstract description of the algorithm is very simple, though a concrete and optimal implementation requires the use of advanced techniques and data structures. The main idea is: i) match all the equivalence rules in the current state of the e-graph, ii) apply the rules creating new e-classes, iii) merge the equivalent e-classes, iv) repeat until saturation (i.e., no changes occur).
@@
The data structure of an e-class stores the information about the e-nodes it contains, a list of the parent e-nodes, and additional information also referred to as semantic analysis. This implementation also maintains a database of patterns that allows the algorithm to efficiently match patterns inside the e-graph structure.
@@
The e-graph is commonly used to represent a single program or expression and their equivalent forms, when applying to simplification. But, the structure can keep any number of expressions as long as we keep a list of the e-classes ids that represents the root of each expression. For example, if we insert the expressions into the e-graph, we would end up with Fig. 1a, minus the equivalent relations. As a result, we would keep the list 222Underlined numbers represent e-class ids. representing the ids of the expressions we inserted. New expressions are added bottom-up. Starting from a terminal, the algorithm checks whether it already exists in the e-graph returning its e-class id if it does, otherwise, it creates a new id. When adding an internal node, the algorithm first converts it to an e-node by replacing its children by their e-class ids and then it checks whether it already exists in the e-graph, returning the corresponding e-class id or a new one. In our example from Fig. 1, if we try to add the expression , it would first retrieve the e-class ids for the terminals and , then it would return the e-class id corresponding to the e-node . Finally, it would create a new e-class with id and the e-node . This mechanism allows us to compactly store a set of expressions and readily assert whether an expression already exists in the structure.
@@
The proposed algorithm, eggp (e-graph genetic programming), follows the same structure as the traditional GP. The initial population of size is created using ramped half-and-half respecting a maximum size and maximum depth (Koza, 1992) and, for a number of generations, it will choose pair of parents using tournament selection, applying the subtree crossover with probability followed by the subtree mutation with probability , replacing the offsprings following a certain criteria. The key differences of eggp are:
a single step of equality saturation is executed after inserting new expressions, merging equivalent expressions.
@@
Notice that a single step of equality saturation will not guarantee the insertion of all equivalent expressions in the e-graph but, if we apply more iterations, the e-graph can grow exponentially large. This issue is amplified by the fact that we are storing multiple expressions. As we will see in Sec. 6, the single step seems to be sufficient to improve the results and the benefits of increasing the number of steps is a subject for future research.
We implemented single and multi-objective versions (called eggpso and eggpmo), with eggpso replacing the current population with the generated offspring and eggpmo replacing it by the set of individuals formed by: the Pareto front, the next front after excluding the first Pareto-front, and a selection of the last offspring at random until it reaches the desired population size. Keeping two ranks of dominance and filling up the remainder of the population with new expressions is meant to stimulate the combination of new expressions (exploration) while keeping the best fronts (exploitation). At the end of the execution, we can extract the Pareto-front from the entire history of the search, this is equivalent to the traditional NSGA-II algorithm (Deb et al., 2000) as the dominance relation is transitive.
Moreover, we keep a database of generated expressions sorted by fitness and size (both objectives used in this work), so the Pareto-front can be retrieved in where is the number of extracted individuals. A new expression can be inserted into this structure in , where is the number of evaluated expressions so far. Finally, if we add the expression , equality saturation will generate the equivalent form (constants are replaced by parameters), storing it in the database of expressions with size .
4.1. E-graph crossover and mutation
The e-graph crossover and mutation operators exploits the information of the search history stored on the e-graph to increase the probability of generating an unvisited expression. The e-graph crossover (Fig. 2a) involves two parents chosen with tournament selection and replaces a random subtree of the first parent with a random subtree sampled from a subset of all possible subtrees of the second parent. This subset is built such that, when replacing the chosen subtree of the first parent, it will generate an unvisited expression. In the event that this set is empty or the algorithm chooses not to perform the crossover (with probability ), it will return the unmodified first parent.
The e-graph mutation (Fig. 2b) is applied to the offspring of the crossover with a probability . Like the traditional subtree mutation, it replaces a random subtree of that solution with a randomly generated subtree using either the grow or full method (chosen at random). After the new expression is created, it is checked whether it already exists in the current e-graph. If it does, the node at the root of the generated subtree is exchanged by another node chosen at random from a subset of the symbols with the same arity. This subset is formed by all the symbols that would create an unvisited expression. When this subset is empty, the current mutated expression is returned.
@@ 4.1. E-graph crossover and mutation
The full implementation of eggp has nine hyperparameters: number of generations, population size, maximum expression size, loss function (MSE, Gaussian, Poisson, Bernoulli, ROXY (Lelli et al., 2017)), number of iterations and retries for the parameter optimization, probabilities of crossover and mutation, and the list of non-terminals.
@@ 5. Experiments
Operon is a carefully crafted implementation of GP for symbolic regression with runtime performance in mind and a good set of default hyperparameters. It incorporates multiple mutation operators that allow a finer perturbation of a solution. Besides, it envelops every variable node with a scaling parameter adjusted using nonlinear optimization. PySR also supports the same mutation operators and nonlinear optimization of the parameters, it stands out with the use of an island model capable of keeping the diversity of the population to stimulate the exploration of the search space. It also applies a simplification heuristic on a selection of the expressions. Both PySR and Operon uses multi-objective optimization with accuracy and expression size as the default objectives.
@@
We have fixed all the common hyperparameters to an empirically set of default values. The only differences in settings are: for PySR we are using populations in its island model, so the size of each population is of the population size for the other algorithms, earlier experiments revealed that PySR performs significantly worse when using a single population; for Operon, we perform a maximum of optimization iterations instead of iterations with different starting points since it does not support multiple restarts; for eggp, of the training data is separated and used as a validation set to calculate the fitness, while the parameters are fitted using the remaining .
@@
| pop. size / gens. / tourn. size | |
@@
For the first set of experiments, we evaluated each algorithm using the recently proposed333https://github.com/cavalab/srbench/discussions/174#discussioncomment-10285133 reduced SRBench benchmark. This set is supposed to be representative as it contains datasets with different characteristics. For this experiment, we applied a -fold cross-validation repeating the experiment times, creating a total of runs. Finally, we picked some real-world datasets from the literature corresponding to data from different fields, these data already have pre-determined training and test sets, as such we will run repetitions of each experiment. For every experiment, we store the final Pareto front and will report the performance plot, the area under the curve (AUC), average rank among all datasets, statistical test of the ranks, the average and standard deviation of the running time. All algorithms were restricted to a single core to ensure equal conditions. Table 2 shows the datasets with their corresponding number of data points, features and the chosen maximum size parameter.